OpenStudy (anonymous):

roberta needs ribbon for a craft project. the ribbon sells for $3.75 per yard. whats the cost in dollars for 48 inches of ribbon ?

OpenStudy (anonymous):

1 yd = 36 inches. 36/3.75 = 48/x (36 inches to 3.75 = 48 inches to x$) cross multiply (36)(x) = (3.75)(48) 36x = 180 x = 180/36 x = 5 If I did this correctly, it would cost $ 5.00. You might need to get a second opinion just to be sure.

OpenStudy (ryan123345):

thats incorrect.

OpenStudy (anonymous):

@ryan123345...where are you getting the 3 from ? The one you got above the x

OpenStudy (anonymous):

I hope you don't think 48 inches is 3 yards

OpenStudy (anonymous):

@amistre64 ......did I do this problem correct or is Ryan correct ?

OpenStudy (amistre64):

roberta needs ribbon for a craft project. the ribbon sells for $3.75 per yard. whats the cost in dollars for 48 inches of ribbon ? inches to yards are not common measures; we have to convert yards to inches. 3 feet in a yard, 12 inches in a foot. 12*3 = 36 inches in a yard 3.75 per yards is also 3.75 per 36 inches. times that by 48 inches \[\frac{$3.75}{36\cancel{in}}*\frac{42\cancel{in}}{1}=$$$\]

OpenStudy (amistre64):

i get about $4.38 for 42 inches

OpenStudy (amistre64):

lol, i read 48 as 42

OpenStudy (amistre64):

$5 for 48 inches is correct

OpenStudy (anonymous):

Thank you....I was thinking I did it correctly, but Ryan kinds threw me off...lol